# LocalSearchBench: Benchmarking Agentic Search in Real-World Local Life Services

### v1

@@ 3.2. LocalRAG Construction
To enable efficient and accurate merchant information retrieval for LRMs high-performance tool calling, we further construct a comprehensive Retrieval-Augmented Generation (RAG) system, which we term LocalRAG, specifically designed as a tool interface for LRMs. Built upon our previously generated merchant dataset, LocalRAG consists of three core components: vector embedding generation, similarity retrieval, and result reranking, as shown in Figure 6.
3.2.1. Vector Embedding Generation
LocalRAG employs a state-of-the-art embedding model to convert merchant information into high-dimensional vector representations. The process begins by transforming structured merchant data into coherent natural language descriptions. Each merchant record undergoes a text formatting pipeline that systematically concatenates key fields including name, category, address, and service descriptions into readable text. The complete merchant profile is then directly encoded and stored as indexed vector chunks in the vector database, ensuring the embeddings capture both explicit attributes and semantic relationships for nuanced similarity matching.
3.2.2. Similarity Retrieval
The generated embeddings are organized in an indexed vector database with cosine similarity-based search capabilities for fast approximate nearest neighbor retrieval, implementing dense retrieval methodology. During query processing, user queries are embedded using the same model and matched against the indexed merchant vector chunks to retrieve the top-N semantically similar merchant candidates.
3.2.3. Result Reranking
To improve retrieval accuracy, LocalRAG first retrieves the top-N merchants from the vector database as a larger candidate set, then applies a specialized reranking model to perform fine-grained relevance scoring between the query and these candidates. The reranking model computes precise relevance scores for each candidate merchant, ultimately selecting the top-K most relevant merchants based on these refined scores to generate the final answer for LRMs.
Figure 6. Workflow of LocalRAG
@@ 3.3. Multi-hop QA Construction
Figure 7. Workflow of Multi-hop QA Construction.
In this subsection, we present the multi-hop QA construction of LocalSearchBench. Building upon the data construction pipeline described in the previous subsections, our multi-hop QA construction pipeline consists of four main stages: seed question collection, question instantiation, answer collection, and data validation. Each of these stages is detailed in the following subsubsections.
@@ 3.3.1. Seed Question Collection
We collect 1,200 single-hop questions from local life service platform M (90%) and enriched sources (10%) across six service categories: dining, lifestyle, shopping, accommodation, healthcare, and tourism. The platform data captures authentic user queries, while enriched data provides coverage for complex scenarios underrepresented in natural logs. Following the intelligence level taxonomy in Table 3, we focus on L3 (composite requirements) and L4 (personalized planning) questions, as they require multi-step reasoning within a static environment, while L5 tasks involve dynamic cross-platform interactions that are difficult to simulate offline.
(a) Intelligence Level Distribution
(b) Example Questions
Figure 8. Intelligence Level Analysis for Collected Questions
@@ 3.3.2. Question Instantiation
We instantiate the collected questions with city-specific information through a two-phase process. First, we construct city profiles for each of the 3 cities (Section 3.1.1), collecting structured information from public resources including landmarks, cultural events, local cuisine, price bands, etc. The collected city data is preprocessed using automated scripts to remove duplicates, normalize formats, and correct inconsistencies, then validated by native annotators to ensure factual accuracy and cultural authenticity.
Second, annotators from platform M manually instantiate questions into concrete multi-hop queries. For each city, approximately 100 questions are selected from the collected 1,200 questions and grounded with city-specific information. To transform single-hop questions into multi-hop queries, annotators enrich questions with 2–4 reasoning segments (2 for 3-hop, 3 for 4-hop, 4 for 5-hop) based on business relevance and logical coherence. These segments include: cross-merchant comparison, schedule planning, chained events, and constraint resolution. We maintain target intelligence-level and hop-count distributions as shown in Figures 9(a) and 9(b), prioritizing underrepresented categories to ensure balanced coverage across service categories and geographic regions (Sections 3.1.1 and 3.1.1).
(a) Intelligence level distribution of QA
(b) Hop count distribution of QA
Figure 9. Distribution of 300 Multi-hop QA
This process generates 300 grounded question instances. For each question, annotators document the hop count, the retrieval target of each hop, and the required search tool (Web Search or LocalRAG) for each hop, enabling downstream validation to enforce hop-by-hop requirements.
@@ 3.3.3. Answer Collection
For each instantiated question, we generate its reference answer through a systematic two-step process, leveraging our pre-constructed LocalRAG described in Section 3.2 along with LLM and human expert validation to ensure comprehensive and accurate ground truth answers.
Step 1: Merchant Retrieval. For each question, LocalRAG retrieves relevant local merchant from our comprehensive business database. LocalRAG maintains an indexed vector database of merchant information, where each merchant record is encoded as high-dimensional vector representations capturing merchant names, categories, addresses, and service descriptions along with their semantic relationships. Using semantic similarity matching combined with geographical proximity analysis, the system retrieves relevant business listings, location metadata, and service information that directly relate to the question context. For complex multi-step questions, LocalRAG performs multiple retrieval operations to gather comprehensive merchant information across different categories and locations, providing the necessary contextual foundation for subsequent reasoning steps.
Step 2: Initial Answer Generation. With the retrieved merchant context from LocalRAG, we utilize state-of-the-art LLMs (GPT-5 (openai2025gpt5) and Claude-4.1-opus (anthropic2025claude41)), assisted by domain experts using collaborative tools, to generate preliminary answers. This process is guided by prompts that emphasize factual accuracy, completeness, and strict adherence to the provided merchant information. The prompts explicitly direct the models to cite specific merchant details and avoid introducing information not present in the retrieved context, thus minimizing hallucination. The resulting question-answer pairs serve as input for the subsequent validation and refinement stage.
3.3.4. Data Validation
To ensure the highest quality and reliability of our benchmark, all preliminary QA pairs undergo a rigorous data validation process beyond the initial generation and refinement described in Section 3.3.3. This section outlines the final quality assurance process designed to meet benchmark-level standards, as illustrated in Figure 7. This multi-phase procedure is conducted by experienced technical specialists with deep domain expertise in local life services and involves both answer revision and quality filtering.
Step 1: Expert Answer Revision. Each generated answer is manually reviewed by our expert team consisting of three paper authors working independently. With the assistance of AI-powered development tools, each reviewer independently verifies the factual accuracy of every claim, assesses the completeness of the explanation, and refines the language for clarity and precision. This human-in-the-loop approach allows for nuanced corrections that automated systems might miss, ensuring each answer is not only correct but also comprehensible and contextually appropriate.
Step 2: Quality Filtering. After revision, the QA pairs are subjected to a final filtering step. QA pairs are retained only if they meet all of the following quality standards: (1) questions are clear, well-formulated, and unambiguous; (2) answers are factually correct and complete after revision; and (3) answers can be sufficiently grounded in real-world local service scenarios and business logic. We also enforce balanced distribution across the service categories and geographic regions defined in Section 3.1.1. This stringent filtering ensures that only the most clear, correct, and valuable QA pairs are included in the final benchmark.
3.3.5. Statistics of Multi-hop QA
Figure 9 summarizes the QA pairs, whose intelligence-level and hop-count distributions align with Figures 9(a) and 9(b).
@@ 4. LocalPlayground
To verify the effectiveness of the proposed benchmark, we design LocalPlayground, an agent evaluation framework that employs autonomous agents to assess LLM performance, as shown in Figure 10. The framework consists of two key components that work collaboratively: a Search Agent that retrieves and synthesizes information, and a Validation Agent that evaluates response quality against ground truth.
Search Agent
The Search Agent retrieves relevant information using two search tools: LocalRAG and Web Search.
LocalRAG performs semantic similarity matching and geographical proximity analysis on vector-indexed merchant databases to retrieve business listings, points of interest and service descriptions.
Web Search queries real-time contextual information such as temporal facts, weather conditions, general knowledge, news and events and supplementary product details that may affect merchant operations and user decisions through web requests to search APIs, which interface with search engines (specifically Baidu Search (liu2021pre) in our experiments) to retrieve up-to-date web content.
The Search Agent operates iteratively over N rounds. In each round, it invokes LocalRAG at most once and Web Search at most once, allowing the agent to progressively refine its information gathering strategy based on intermediate results.
Validation Agent
To ensure comprehensive and objective assessment of generated outputs, we employ a Validation Agent implemented as an LLM-as-a-Judge framework, as shown in Figure 10. Given the generated outputs and the golden answer, an LLM (specifically Claude-Sonnet-4 (anthropic2025claude4) in our experiments) evaluates answer quality across seven dimensions: (1) Correctness: measures whether the answer is factually accurate; (2) Completeness: measures whether the answer fully addresses the question, (3) Fluency: measures the readability and coherence of the response; (4) Faithfulness: measures the degree to which the response avoids fabricated or hallucinated content; (5) Safety: measures whether the response avoids harmful or inappropriate content; (6) Avg. tool calls: measures the average number of tool invocations (LocalRAG and web search) and (7) Avg. rounds: measures the average number of conversation rounds.
Dimension (1) employs a binary scoring scheme (0 or 1), clearly distinguishing correct from incorrect responses. For dimensions (2)-(5), each dimension is evaluated using a granular 10-point rating system (0-10), with higher scores indicating superior performance. Scores across dimensions (1)-(5) are normalized and expressed as percentages for standardized comparison. For dimensions Dimensions (6)-(7) capture continuous variables representing frequency counts. To ensure reliability, each instance is evaluated five times per dimension, with final scores computed by averaging across trials. Candidate systems are anonymized, question order is randomized, and the judge model is separate from all evaluated models to prevent self-assessment.
Figure 10. Overview of LocalPlayground
@@ 5. Evaluation
We aim to answer the following research questions:
RQ1 (Benchmark Quality): How well does LocalSearchBench capture the diversity and complexity of real-world merchant search scenarios?
RQ2 (Benchmark Results): How do existing open-source and closed-source LLMs perform on our benchmark?
RQ3 (Sensitivity Study): How does the maximum number of conversation rounds affect search agent performance?
@@ 5.1. Evaluated Models
We evaluate ten widely recognized LRMs using LocalPlayground, including five closed-source LRMs, i.e., GPT-4.1 (openai2025gpt4-1), Gemini-2.5-Pro (comanici2025gemini), Qwen-Plus, Hunyuan-T1 (team2025hunyuan), and LongCat-Large-32K (team2025longcat) as well as five open-source LRMs, i.e., DeepSeek-V3.1 (deepseek2025v3), GLM-4.5 (zeng2025glm), Qwen3-235B-A22B, Qwen3-32B, and Qwen3-14B (yang2025qwen3).
Unless specified, LocalPlayground is configured with default settings: (1) a maximum of 5 (N = 5) reasoning-action iterations per question; (2) for LocalRAG, Qwen3-Embedding-8B (zhang2025qwen3) as the embedding model and Qwen3-Reranker-8B (zhang2025qwen3) as the reranking model, with dense retrieval returning top-N = 100 candidates and reranking returning top-K = 20 merchants; (3) for web search, top-k retrieval where k = 20; and (4) LRMs decoding temperature set to 0 to ensure reproducibility. All experiments were conducted on a system equipped with an Intel Xeon Gold 5218 CPU @ 2.30GHz and 1 $\times$ H20-141G GPU.
5.2. Benchmark Quality (RQ1)
To evaluate the quality of our LocalSearchBench database (Section 3.1.1), we employ the two-stage validation framework described in Section 3.1.3. We randomly sample 1,000 merchant entries and adopt an LLM-as-judge approach using Claude-Sonnet-4 (anthropic2025claude4) to assess data quality across multiple dimensions. As shown in Table 4, both augmented and privacy rewriting data demonstrate high quality with overall scores of 0.8596 and 0.9217, respectively.
To further validate the practical utility of our benchmark and mitigate potential biases inherent in automated evaluation, we invited four independent evaluators with expertise in local life service domains, who each evaluated outputs from all methods to ensure comprehensive coverage. They assessed outputs in a blind manner based on data consistency, completeness, and anonymization effectiveness using a 5-point Likert scale (5-Strongly Agree to 1-Strongly Disagree). For analysis, ratings of 4-5 were converted to 1 (satisfied) and 1-3 to 0 (not satisfied).
As shown in Table 7, the human evaluation achieves 86.51% agreement with automated LLM assessment across all 20 fields, validating the effectiveness of our LLM-as-judge approach.
To ensure the quality of our Multi-hop QA, we applied stringent manual verification (Section 3.3.4), retaining only questions with clear, answerable information from the given merchant data and filtering out ambiguous or unanswerable queries.
Finding 1: Automated LLM evaluation shows high quality across LocalSearchBench, and human verification on sampled data confirms the reliability of automated assessment, validating the benchmark’s practical utility.
Table 4. Evaluation Frameworks for Data Quality Assessment
(a) Augmented Data Quality
Dimension Weight Score
Data Completeness 20% 0.8605
Logical Consistency 30% 0.9000
Geographic Accuracy 10% 0.8568
Content Quality 20% 0.8521
Format Standardization 10% 0.9060
Innovation Level 10% 0.7080
Overall 100% 0.8596
(b) Privacy Rewriting Data Quality
Dimension Weight Score
Anonymization Level 50% 0.9473
Logical Consistency 20% 0.8383
Content Consistency 20% 0.9919
Innovation Level 10% 0.8200
Overall 100% 0.9217
5.3. Benchmark Results (RQ2)
Impact of web search.
We first analyze the impact of web search, comparing LocalPlayground (LRM+ LocalRAG +web search) against LRM + LocalRAG without websearch as the baseline to assess web search effectiveness. We exclude configurations of the base LRM alone or LRM without LocalRAG from our comparison, as multi-hop questions in our benchmark require information exclusively available in the LocalRAG knowledge base.
As shown in Table 5, incorporating web search into LocalPlayground presents a trade-off between answer quality and faithfulness. On average, LocalPlayground improves Correctness by 4.37pp (from 25.58% to 29.95%) and Completeness by 3.95pp (from 73.38% to 77.33%), but reduces Faithfulness by 3.64pp (from 65.63% to 61.99%). This trade-off is most pronounced for DeepSeek-V3.1, which gains 4.57pp in Correctness but loses 8.45pp in Faithfulness. These results suggest that while web search is essential for retrieving critical information in multi-hop reasoning, it also introduces noise that compromises response faithfulness.
Performance across language models.

### v2

@@ 3.2. LocalRAG Construction
To support efficient and accurate merchant retrieval for LRMs, we build LocalRAG, a Retrieval-Augmented Generation system on our merchant database (Figure 8). LocalRAG uses a embedding model to convert each merchant profile into a high-dimensional vector. Structured fields are directly encoded into vectors and indexed in a city-partitioned vector database. During query processing, the target city is first extracted from the query text, then the corresponding city-partitioned vector database is selected. The query is embedded with the LocalRAG embedding model and matched via cosine similarity against merchant vectors from that city, selecting the top-N semantically related candidates through semantic retrieval. LocalRAG then applies a reranking model to the top-N candidates, scoring query-merchant relevance and returning the top-K merchants as the final retrieval results for LLM agents.
@@ 3.3. Multi-hop QA Construction
Figure 6. How Multi-hop QA pairs in LocalSearchBench are constructed.
In this subsection, we present the multi-hop QA construction of LocalSearchBench. Building upon the data construction pipeline described in the previous subsections, our multi-hop QA construction pipeline consists of four main stages: seed question collection, question instantiation, answer collection, and QA validation. Each of these stages is detailed in the following subsubsections.
@@ 3.3.1. Seed Question Collection
We collect 1,200 single-hop questions from local life service platform M (90%) and enriched sources (10%) across 6 service categories (in § 3.1.1). The platform data captures authentic user queries, while enriched data provides coverage for complex scenarios underrepresented in natural logs. Following the intelligence level taxonomy in Table 5, we focus on L3 (Intelligent) and L4 (Expert) questions, as they require multi-step reasoning within a static environment, while L5 (AGI) questions involve dynamic cross-platform interactions that are difficult to simulate offline.
(a) Intelligence Level Distribution
(b) Example Questions
Figure 7. Intelligence Level Analysis for Collected Questions
@@ 3.3.2. Question Instantiation
We instantiate the collected questions with city-specific information through a two-phase process. First, we construct city profiles for each of the 9 cities (§ 3.1.1), collecting structured information from public resources including landmarks, cultural events, local cuisine, price bands, etc. To facilitate consistent evaluation, temporal elements are anchored to static timestamps, creating a fixed reference point for all queries. The collected city data is preprocessed using automated scripts to remove duplicates, normalize formats, and correct inconsistencies, then validated by annotators to ensure factual accuracy and cultural authenticity. Second, annotators from platform M manually instantiate questions into concrete multi-hop queries. For each city, approximately 100 questions are selected from the collected 1,200 questions and grounded with city-specific information. To transform single-hop questions into multi-hop queries, annotators enrich questions with 2–4 reasoning segments based on business relevance and logical coherence. These segments include: cross-merchant comparison, schedule planning, chained events, and constraint resolution. We maintain target intelligence-level and hop-count distributions (Figure 6), prioritizing underrepresented categories to ensure balanced coverage across service categories and cities in § 3.1.1.
@@ 3.3.3. Answer Collection
Figure 8. Overview of LocalPlayground.
For each instantiated question, we generate reference answers through a multi-step process using LocalRAG (§ 3.2), web search, and LLM-based generation. We first check whether the question needs real-time information (e.g., current events, weather, news) or can be answered with merchant data alone. If real-time signals are present, we invoke web search; otherwise, the query relies on the city-partitioned merchant vector database. For all questions, LocalRAG retrieves merchants scoped to the target city. For multi-hop questions, the hop count, retrieval target, and required search tool for each hop are predetermined during question construction. Search tools then execute iterative retrieval per hop, gathering merchants across categories/locations and combining them with any web-search snippets to form the full reasoning context. Claude-4.5-opus (Anthropic, 2025) then generate preliminary answers with expert oversight. The LLM reranks each hop’s retrieved candidates and selects the top-$k$ passages (k=5k{=}5) as evidence, then synthesizes a final answer grounded in these snippets while following the fixed hop plan and the original query. The prompt enforces: (1) grounding in retrieved evidence, (2) alignment with the fixed hop plan, (3) explicit citation of merchant and web evidence, and (4) avoidance of unsupported content. The resulting question-answer pairs serve as input for validation.
3.3.4. QA Validation
All preliminary QA pairs undergo a two-stage validation process (Figure 6). First, hop-by-hop validation ensures: (1) every hop has a clear answer, meaning retrieval targets must be found through LocalRAG or web search with sufficient evidence; (2) strict adherence to the target hop-count chain; (3) conformity to local life service business logic. Questions failing these criteria are manually modified and re-submitted to answer collection (§ 3.3.3). Second, three domain experts independently review answers for factual accuracy, completeness, and clarity, resolving disagreements through consensus. Answers are deterministic with balanced distribution across service categories and cities (§ 3.1.1), yielding 900 high-quality QA pairs for the final benchmark (Figure 6).
@@ 4. LocalPlayground
We design LocalPlayground for automated testing with LocalSearchBench. It contains a Search Agent for multi-hop retrieval and answering, and a Validation Agent for quality examination of answer and reasoning trajectory (Figure 8).
Search Agent. The Search Agent uses two tools: LocalRAG for vector-DB merchant retrieval and web search for real-time information (facts, weather, news, events) via Baidu APIs (Liu et al., 2021). It performs iterative search up to N rounds, calling each tool at most once per round to refine answers. For each query $q_{0}$, it records the trajectory $\tau=\{s_{0},a_{0},s_{1},a_{1},\dots,s_{t}\}$, with states $s_{i}=(q_{i},C_{i})$ storing the current sub-question and accumulated evidence. At step $i$, it selects a reasoning action $a_{i}^{\text{reason}}$ and, if needed, a search action $a_{i}^{\text{search}}$ to obtain evidence $E_{i+1}$, update $C_{i+1}=C_{i}\cup E_{i+1}$, and generate $q_{i+1}$ from $(q_{i},C_{i+1})$. A final reasoning step synthesizes $q_{t}$, and $s_{i+1}=F(s_{i},a_{i})$ forms a directed chain over sub-questions and evidence for reasoning.
Validation Agent. We employ an LLM-as-judge (Claude-4.5-opus) to evaluate each agent interaction along two dimensions: answer quality and trajectory quality. Answer quality is assessed on 7 metrics: correctness (factual accuracy), completeness (covers the query), fluency (readability), faithfulness (no hallucination), safety (no harmful content), avg. tool calls and avg. rounds. Each query is associated with a multi-attribute constraint checklist defining all necessary and sufficient conditions for a valid answer. During evaluation, the Validation Agent verifies whether the agent’s retrieved merchant satisfies every constraint on the checklist. If multiple merchants meet all listed conditions, each is considered a correct answer. Trajectory quality is scored on 4 dimensions: action relevance (step aligns with intent), evidence sufficiency (evidence is enough), causal coherence (later steps correctly reference earlier evidence), and efficiency (avoids redundant/ineffective calls).
Each instance is evaluated five times per dimension and averaged. Systems are anonymized, question order is randomized, and the judge model is distinct from all evaluated models.
Table 6. Performance comparison of different models on LocalSearchBench (Max Round N = 5). The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
Models Answer Evaluation Metrics Trajectory Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
Qwen3-235B-A22B (w/o thinking) 2.00 2.93 21.20 50.94 69.16 25.28 79.72 75.22 43.61 50.68 45.99
Qwen3-235B-A22B (w/ thinking) 2.31 3.17 30.20 71.20 71.58 26.90 81.76 80.42 45.75 52.04 48.63
Qwen3-32B (w/o thinking) 2.78 3.11 19.80 40.96 68.50 21.38 80.76 74.67 46.52 48.82 49.84
Qwen3-32B (w/ thinking) 2.80 3.12 25.60 40.66 68.44 22.40 79.54 75.06 46.99 48.87 49.13
Qwen3-14B (w/o thinking) 2.53 2.07 24.20 40.60 69.62 27.44 80.78 80.46 46.44 50.96 50.02
Qwen3-14B (w/ thinking) 2.57 2.12 25.20 40.98 69.32 28.40 80.44 81.19 47.24 52.52 51.65
GPT-4.1 1.73 2.42 18.50 45.37 65.93 28.76 77.38 68.47 38.62 45.83 42.29
o3(high) 2.91 3.38 31.50 69.64 70.72 33.89 81.87 75.93 44.71 51.78 42.96
Gemini-2.5-Flash 1.84 2.51 21.00 58.44 68.04 35.72 79.70 70.58 41.61 48.94 46.91
Gemini-2.5-Pro 2.75 3.10 32.30 71.03 71.12 34.93 82.32 77.31 45.73 52.87 41.78
LongCat-Flash-Chat 2.34 3.07 25.30 52.98 69.45 27.49 83.61 77.78 47.33 50.86 52.29
LongCat-Flash-Thinking 3.04 3.20 30.70 68.83 69.07 31.47 80.10 78.50 47.37 53.18 53.27
GLM-4.6 (w/o thinking) 2.86 3.86 29.00 76.45 70.37 35.40 81.40 74.28 48.44 50.79 52.76
GLM-4.6 (w/ thinking) 3.08 4.06 32.80 76.83 70.27 37.48 81.30 77.66 48.90 52.67 54.43
Deepseek-V3.2 (w/o thinking) 3.19 4.18 31.70 77.08 70.12 35.96 81.52 75.51 48.49 52.23 54.33
Deepseek-V3.2 (w/ thinking) 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
Table 7. Ablation study of tool integration on thinking models using LocalSearchBench (Max Round N = 5). The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
Methods Answer Evaluation Metrics Trajectory Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
LocalRAG workflow 1.00 1.00 15.18 58.30 70.80 66.10 85.40 37.70 23.53 32.77 39.35
Gemini-2.5-Pro (w/o tools) 0.00 1.00 0.00* 38.42 91.65 9.82 99.89 25.21 8.72 21.70 13.62
+ LocalRAG 1.86 2.83 23.11 63.92 82.32 38.65 84.95 68.10 37.37 46.93 39.96
+ LocalRAG & web search 2.75 3.10 32.30 71.03 71.12 34.93 82.32 77.31 45.73 52.87 41.78
GLM-4.6 (w/o tools) 0.00 1.00 0.00* 36.06 91.92 12.12 99.67 46.88 7.92 28.85 17.08
+ LocalRAG 2.33 3.32 23.91 69.21 83.75 41.45 84.72 63.33 36.21 48.02 52.05
+ LocalRAG & web search 3.08 4.06 32.80 76.83 70.27 37.48 81.30 77.66 48.90 52.67 54.43
Deepseek-V3.2 (w/o tools) 0.00 1.00 0.00* 39.30 90.20 18.56 99.55 52.65 10.80 32.10 41.15
+ LocalRAG 2.88 3.87 27.04 70.05 82.98 40.59 83.85 65.55 42.78 47.85 54.07
+ LocalRAG & web search 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
* The retrieved merchant must have a poi_id from LocalRAG.
Table 8. Sensitivity analysis of Max Round N on model performance. The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
Max Round Answer Evaluation Metrics Trajectory Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
$N=2$ 2.46 2.82 30.50 70.61 76.44 54.65 89.94 59.94 31.17 44.82 40.77
$N=3$ 2.78 3.30 33.50 74.79 73.57 43.50 84.58 61.68 42.96 47.42 48.50
$N=4$ 3.03 3.95 34.20 76.34 72.27 41.26 82.53 72.97 49.28 51.32 54.13
$N=5$ 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
$N=6$ 3.76 4.75 34.90 74.52 70.36 38.56 80.91 68.63 49.28 51.65 47.10
$N=7$ 3.87 4.85 33.70 73.60 70.19 37.36 80.84 69.50 49.37 48.20 44.99
$N=8$ 4.02 5.01 32.40 73.15 70.24 37.35 80.39 66.89 49.39 47.94 43.58
@@ 5.1. Evaluated Models
We evaluate 16 state-of-the-art thinking and non-thinking models for agents: GPT series (GPT-4.1, o3 (high)) (OpenAI, 2025), Gemini series (Gemini-2.5-Flash, Gemini-2.5-Pro) (Comanici et al., 2025), LongCat series (LongCat-Flash-Chat, LongCat-Flash-Thinking) (Team et al., 2025), Deepseek-V3.2 (DeepSeek-AI, 2025), GLM-4.6 (Zeng et al., 2025), and Qwen3 series (Qwen3-14B, Qwen3-32B, Qwen3-235B-A22B) (Yang et al., 2025a). Unless specified, LocalPlayground uses default settings: maximum tool call rounds N = 5; LocalRAG with Qwen3-Embedding-8B and Qwen3-Reranker-8B (Zhang et al., 2025c) (top-100 dense retrieval, top-20 reranking); web search top-k = 20; and LLM temperature = 0. Experiments run on Intel Xeon Gold 5218 @ 2.30GHz with 1 $\times$ H20-141G GPU.
5.2. Data and LocalRAG Quality
We evaluate LocalSearchBench database with the two-stage framework in §3.1.3, sampling 1,000 merchants. As shown in Table 18, the augmented and privacy-rewritten data achieve high overall quality scores of 0.86 and 0.92 (out of 1).
[…]
Since retrieval quality directly affects downstream QA performance, we evaluate LocalRAG on 27,360 hop-level queries using NDCG@10/20 and MRR@10/20 over the top-20 results of the full retrieval pipeline (§3.2). Our best configuration, Qwen3-Embedding-8B, reaches NDCG@10/20 of 0.84/0.87 and MRR@10/20 of 0.82/0.83 (Table 9). These hop-level queries come from the multi-hop QA construction in §3.3 with ground-truth evidence for each hop; metric definitions and full retrieval results are detailed in Appendix A.1.
5.3. Overall Performance